# Spatially-Grounded Document Retrieval via Patch-to-Region Relevance Propagation

### v1

@@ Abstract
Vision-language models (VLMs) like ColPali achieve state-of-the-art document retrieval by embedding pages as images and computing fine-grained similarity between query tokens and visual patches. However, they return entire pages rather than specific regions, limiting utility for retrieval-augmented generation (RAG) where precise context is paramount. Conversely, OCR-based systems extract structured text with bounding box coordinates but lack semantic grounding for relevance assessment. We propose a hybrid architecture that unifies these paradigms: using ColPali’s patch-level similarity scores as spatial relevance filters over OCR-extracted regions. We formalize the coordinate mapping between vision transformer patch grids and OCR bounding boxes, introduce intersection metrics for relevance propagation, and establish theoretical bounds on retrieval precision. Our approach operates at inference time without additional training. We release Snappy, an open-source implementation demonstrating practical applicability, with empirical evaluation ongoing.
@@ 1 Introduction
Vision-language models (VLMs) such as ColPali (Faysse et al., 2025) have achieved state-of-the-art performance on document retrieval benchmarks by embedding document pages directly as images. ColPali produces 1,024 patch embeddings (a $32\times 32$ grid) per page, each projected to 128 dimensions. The model computes relevance through late interaction, specifically a MaxSim operation that sums the maximum similarity between each query token and all document patches. This approach elegantly sidesteps OCR errors and preserves layout semantics. However, ColPali and its variants return entire pages as retrieval units. For RAG applications, this is problematic: feeding a full page into a language model’s context window introduces irrelevant content, increases latency, inflates costs, and, critically, dilutes the signal that the model must attend to. The retrieval system knows which page contains the answer but not where on the page.
@@
We observe that these paradigms are complementary. ColPali’s patch-level similarity scores encode where on a page the model attends when processing a query. This information is computed but discarded when returning page-level results. OCR systems know what content exists and where it is located, but not why it matters. By unifying these signals through spatial coordinate mapping, we achieve region-level retrieval: returning only the document regions that are both structurally coherent (via OCR) and semantically relevant (via VLM attention).
@@ 1.1 Contributions
    Coordinate Mapping Formalism. We formalize the mathematical correspondence between vision transformer patch grids and OCR bounding boxes, enabling spatial alignment between heterogeneous representations (Section 3.2).
@@
    Relevance Propagation via Interpretability Maps. We repurpose ColPali’s late interaction mechanism to generate per-query-token similarity heatmaps, then propagate these scores to OCR regions through IoU-weighted patch-region intersection (Section 3.3).
@@
    Two-Stage Retrieval Architecture. We introduce a mean-pooling strategy that compresses patch embeddings along spatial axes, enabling efficient candidate retrieval before full-resolution reranking (Section 3.4).
@@
    Theoretical Analysis. We establish bounds on localization precision as a function of patch resolution, derive expected context reduction and precision amplification factors, and analyze computational complexity tradeoffs (Section 4).
@@
    Open Implementation. We release Snappy, a complete system implementing this architecture with ColPali, DeepSeek OCR, and Qdrant vector search, demonstrating practical applicability (Section 5).
We hypothesize that region-level retrieval will improve downstream RAG answer quality while reducing context length, validating the premise that retrieval granularity directly impacts generation fidelity. Empirical evaluation is ongoing and will be reported in a subsequent revision.
@@ 2 Background and Related Work
2.1 Vision-Language Document Retrieval
@@
The ViDoRe benchmark (Faysse et al., 2025) evaluates visual document retrieval across diverse domains, measuring NDCG@5 for page-level retrieval. ColPali achieves strong performance, but the benchmark, like the model, operates at page granularity, leaving region-level retrieval unexplored.
@@ 2.2 Layout-Aware Document Understanding
The LayoutLM family (Xu et al., 2020, 2021; Huang et al., 2022) pioneered joint modeling of text, layout, and visual features for document understanding. LayoutLMv3 introduced Word-Patch Alignment (WPA) pre-training, which predicts whether image patches corresponding to text words are masked. While conceptually related to our patch-OCR alignment, WPA operates at pre-training time to improve representations, whereas our approach uses patch similarities at inference time for retrieval filtering. Critically, LayoutLM models are designed for document understanding tasks (NER, classification) rather than retrieval—they lack late interaction mechanisms and query-conditioned relevance scoring.
OCR-free approaches including Donut (Kim et al., 2022) and Pix2Struct (Lee et al., 2023) perform document understanding directly from pixels. UDoP (Tang et al., 2023) unifies vision, text, and layout modalities. These models excel at understanding but do not address the retrieval problem we target.
@@ 2.3 Region-Level Document Retrieval
RegionRAG. The closest existing work is RegionRAG (Li et al., 2025), which shifts retrieval from document-level to semantic region-level granularity. However, RegionRAG uses a hybrid approach combining bounding box annotations with weakly-supervised signals from unlabeled data and a region-level contrastive loss. Our approach is fundamentally different: we achieve region-level retrieval at inference time using ColPali’s emergent patch attention, requiring no additional training or annotation.
DocVLM. DocVLM (Shpigel Nacson et al., 2024) integrates OCR into VLMs by compressing OCR features into learned queries. This represents the opposite direction of our approach. DocVLM adds OCR to enhance VLM understanding, whereas we use VLM patches to filter and score OCR output for retrieval.
Table 1 summarizes the positioning of our approach relative to prior work.
Table 1: Comparison with related approaches. Our approach is unique in achieving region-level retrieval at inference time without additional training, by propagating VLM patch similarities to OCR bounding boxes.
@@ 3.1 Problem Formulation
Let $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ denote the set of pages in the corpus, where each page $p_{i}$ has an associated set of OCR regions $\mathcal{R}(p_{i})=\{r_{1},\ldots,r_{m}\}$. Each region $r_{j}$ is characterized by its bounding box $B(r_{j})=(x_{1},y_{1},x_{2},y_{2})$ in pixel coordinates and its text content $T(r_{j})$. Our goal is to compute a relevance score $\text{rel}(q,r_{j})$ for each region that captures both semantic relevance and spatial grounding.
@@ 3.2 Coordinate Mapping: Patches to Bounding Boxes
ColPali encodes each page as a grid of $G\times G$ patch embeddings ($G=32$ for ColPali-v1.3) over an input image of resolution $I\times I$ ($I=448$). Each patch corresponds to an $s\times s$ pixel region where $s=I/G=14$ pixels. Patches are indexed in raster scan order (left-to-right, top-to-bottom).
@@ 3.3 Relevance Propagation via Patch Similarity
For OCR region $r$ with scaled bounding box $B^{\prime}(r)$, we propagate patch scores via IoU-weighted aggregation:
$\text{rel}(q,r)=\sum_{j}\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))\cdot\text{score}_{\text{patch}}(j)$ (6)
where the sum is over all patches $j$ with non-zero intersection. This weights each patch’s contribution by its spatial overlap with the region, ensuring that patches fully contained within the region contribute more than peripheral patches.
@@ 3.4 Two-Stage Retrieval Architecture
Stage 1: Candidate Retrieval. We compress patch embeddings via mean pooling along spatial axes to obtain a single page-level embedding. This enables efficient approximate nearest neighbor search (via Qdrant) to retrieve top-$K$ candidate pages. Note that pooling discards spatial information, which may impact recall for pages with small relevant regions (see Section 7 for discussion of this limitation).
Stage 2: Region Reranking. For each candidate page, we compute full patch-level similarity and propagate scores to OCR regions as described in Section 3.3. Regions are ranked by their relevance scores, and top-$k$ regions are returned.
@@ 3.5 Aggregation Strategies
We consider alternative aggregation strategies for propagating patch scores to regions:
@@
$\text{rel}_{\text{IoU}}(q,r)=\sum_{j}\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))\cdot\text{score}_{\text{patch}}(j)$ (9)
The choice of aggregation strategy may affect retrieval quality depending on region size and content density. We plan to empirically compare these strategies in our evaluation.
@@ 4 Theoretical Analysis
4.1 Precision Bounds
The spatial precision of our approach is fundamentally bounded by patch resolution. We formalize this tradeoff.
Theorem 1 (Localization Precision Bound).
For an OCR region with bounding box of width $w$ and height $h$ (in model coordinates), and patch size $s$, the maximum achievable localization precision is:
$\text{precision}\leq\frac{w\cdot h}{(w+s)\cdot(h+s)}$ (10)
@@ 4.1 Precision Bounds
Consider a region with bounding box $B$ of dimensions $w\times h$. The set of patches intersecting $B$ depends on $B$’s alignment with the patch grid. In the worst case, $B$ is positioned such that it intersects partial patches on all four edges. Let the region’s top-left corner fall at position $(x,y)$ within a patch. The region then spans from patch column $\lfloor x/s\rfloor$ to $\lfloor(x+w)/s\rfloor$ and from patch row $\lfloor y/s\rfloor$ to $\lfloor(y+h)/s\rfloor$. The number of intersecting patches is at most $\lceil w/s+1\rceil\cdot\lceil h/s+1\rceil$. The total area covered by these patches is at most $(w+s)\cdot(h+s)$, achieved when the region is maximally misaligned with patch boundaries. Since the region’s area is $w\cdot h$, the precision (ratio of relevant area to retrieved area) is bounded by $(w\cdot h)/((w+s)\cdot(h+s))$. ∎
@@
  • A typical paragraph region ($200\times 50$ pixels): precision $\leq 73\%$
  • A table cell ($100\times 30$ pixels): precision $\leq 60\%$
  • A small label ($50\times 20$ pixels): precision $\leq 46\%$
@@ 4.3 Expected Performance Bounds
Beyond spatial precision, we analyze the expected improvements in retrieval quality and downstream efficiency. These bounds provide theoretical justification for our approach pending empirical validation.
@@ 4.3.1 Context Reduction
4.3.2 Precision Amplification
We model retrieval precision as the probability that a returned region contains relevant content.
Definition 3 (Region Relevance).
Let $R_{j}\in\{0,1\}$ indicate whether region $j$ contains content relevant to query $q$. For a typical factoid query, we assume $\sum_{j}R_{j}\in\{1,2\}$ (one or two relevant regions per page).
Theorem 3 (Precision Comparison).
Let $\rho=\sum_{j}R_{j}/M$ denote the fraction of relevant regions on a page. Then:
(i) OCR-only (random selection of $k$ regions):
$\mathbb{E}[\text{Precision}_{\text{OCR}}]=\rho$ (12)
(ii) ColPali page-level:
$\text{Precision}_{\text{ColPali}}=\rho\quad\text{(all regions returned implicitly)}$ (13)
(iii) Hybrid with score-based ranking: If patch scores $s_{j}$ satisfy $\mathbb{E}[s_{j}|R_{j}=1]>\mathbb{E}[s_{j}|R_{j}=0]$ (relevant regions score higher on average), then selecting top-$k$ by score yields:
$\mathbb{E}[\text{Precision}_{\text{Hybrid}}]>\rho$ (14)
with the improvement depending on the score separation between relevant and irrelevant regions.
Proof.
Parts (i) and (ii) follow directly from the definition of $\rho$. For (iii), score-based ranking is equivalent to a classifier with ROC curve above the diagonal when relevant regions have higher expected scores. By the Neyman-Pearson lemma, thresholding on any statistic positively correlated with relevance improves precision over random selection. ∎
4.3.3 Signal-to-Noise Ratio
We formalize the intuition that region-level retrieval improves the “signal” available to downstream LLMs.
Definition 4 (Retrieval SNR).
For retrieved context $C$, define:
$\text{SNR}(C)=\frac{|C\cap\text{Relevant}|}{|C\cap\text{Irrelevant}|}$ (15)
where $|\cdot|$ denotes token count.
Theorem 4 (SNR Improvement).
Under the assumption that relevant and irrelevant regions have similar average token counts:
(i) Page-level retrieval: $\text{SNR}_{\text{page}}=\rho/(1-\rho)$
(ii) Hybrid top-$k$ retrieval with precision $P_{k}$: $\text{SNR}_{\text{hybrid}}=P_{k}/(1-P_{k})$
The SNR improvement factor is:
$\frac{\text{SNR}_{\text{hybrid}}}{\text{SNR}_{\text{page}}}=\frac{P_{k}(1-\rho)}{\rho(1-P_{k})}$ (16)
Proof.
For page-level retrieval, the fraction of relevant tokens equals $\rho$ by definition. For hybrid retrieval with precision $P_{k}$, the expected fraction of relevant tokens among returned regions is $P_{k}$. The SNR ratio follows from algebraic manipulation. ∎
4.3.4 Summary of Expected Gains
Table 2 provides a unified view of the cost-quality tradeoff, showing that our hybrid approach achieves the context efficiency of OCR-based selection while providing the relevance ranking that OCR alone cannot offer.
Table 2: Combined efficiency-quality comparison. The hybrid approach uniquely achieves both low context cost and high precision.
Method Context Cost Precision Best Use Case
[…]

### v2

@@ Abstract
Late-interaction multimodal retrieval models like ColPali achieve state-of-the-art document retrieval by embedding pages as images and computing fine-grained similarity between query tokens and visual patches. However, they return entire pages rather than specific regions, limiting utility for retrieval-augmented generation (RAG) where precise context is paramount. Conversely, OCR-based systems extract structured text with bounding box coordinates but lack semantic grounding for relevance assessment. We propose a hybrid architecture that unifies these paradigms: using ColPali’s patch-level similarity scores as spatial relevance filters over OCR-extracted regions. We formalize the coordinate mapping between vision transformer patch grids and OCR bounding boxes, introduce intersection metrics for relevance propagation, and establish theoretical bounds on area efficiency. We evaluate on BBox-DocVQA with ground-truth bounding boxes. For within-page localization (given correct page retrieval), ColQwen3-4B with percentile-50 thresholding achieves 59.7% hit rate at IoU@0.5 (84.4% at IoU@0.25, 35.8% at IoU@0.7), with mean IoU of 0.569, compared to $\sim$6.7% for random region selection. Our approach reduces context tokens by 28.8% compared to returning all OCR regions and by 52.3% compared to full-page image tokens. Our approach operates at inference time without additional training. We release Snappy, an open-source implementation at https://github.com/athrael-soju/Snappy.
@@ 1 Introduction
Late-interaction retrievers such as ColPali (Faysse et al., 2025) have achieved state-of-the-art performance on document retrieval benchmarks by embedding document pages directly as images. ColPali produces 1,024 patch embeddings (a $32\times 32$ grid) per page, each projected to 128 dimensions. The model computes relevance through late interaction, specifically a MaxSim operation that sums the maximum similarity between each query token and all document patches. This approach elegantly sidesteps OCR errors and preserves layout semantics. However, ColPali and its variants return entire pages as retrieval units. For RAG applications, this is problematic: feeding a full page into a language model’s context window introduces irrelevant content, increases latency, inflates costs, and, critically, dilutes the signal that the model must attend to. The retrieval system knows which page contains the answer but not where on the page.
@@
We observe that these paradigms are complementary. ColPali’s patch-level similarity scores encode where on a page the model attends when processing a query. This information is computed but discarded when returning page-level results. OCR systems know what content exists and where it is located, but not why it matters. By unifying these signals through spatial coordinate mapping, we achieve region-level retrieval: returning only the document regions that are both structurally coherent (via OCR) and semantically relevant (via late-interaction attention).
@@ 1.1 Contributions
    Coordinate Mapping Formalism. We formalize the mathematical correspondence between vision transformer patch grids and OCR bounding boxes, enabling spatial alignment between heterogeneous representations (Section˜3.2).
@@
    Relevance Propagation via Interpretability Maps. We repurpose ColPali’s late interaction mechanism to generate per-query-token similarity heatmaps, then propagate these scores to OCR regions through IoU-weighted patch-region intersection (Section˜3.3).
@@
    Two-Stage Retrieval Architecture. We introduce a two-stage architecture that enables efficient candidate retrieval before full-resolution region-level reranking (Section˜3.4).
@@
    Theoretical Analysis. We establish bounds on area efficiency as a function of patch resolution, derive expected context reduction factors, and analyze computational complexity tradeoffs (Section˜4).
@@
    Empirical Validation and Open Implementation. We evaluate on BBox-DocVQA, demonstrating 59.7% hit rate at IoU@0.5 with 52.3% token savings versus full-page retrieval (Section˜6). We release Snappy, a complete open-source system implementing this architecture (Section˜5).
  6. 6.
    Model and Training Data Analysis. We analyze how model architecture and training data affect patch-to-region localization, identifying two distinct regimes: a model-dependent regime where larger models yield substantial gains, and a precision-bound regime where patch quantization limits dominate regardless of model capacity (Section˜6.5).
On BBox-DocVQA, 59.7% of retrieved regions achieve IoU $\geq$ 0.5 with ground-truth evidence bounding boxes, while reducing context tokens by 52.3% compared to full-page retrieval.
@@ 2 Background and Related Work
2.1 Late-Interaction Multimodal Retrieval
@@
The ViDoRe benchmark (Faysse et al., 2025) evaluates visual document retrieval across diverse domains, measuring NDCG@5 for page-level retrieval. ColPali achieves strong performance, but the benchmark, like the model, operates at page granularity; region-level retrieval remains unexplored.
@@ 2.2 Layout-Aware Document Understanding
The LayoutLM family (Xu et al., 2020, 2021; Huang et al., 2022) pioneered joint modeling of text, layout, and visual features for document understanding. LayoutLMv3 introduced Word-Patch Alignment (WPA) pre-training, which predicts whether image patches corresponding to text words are masked. DocFormer (Appalaraju et al., 2021) extends this direction with multi-modal self-attention that correlates text and visual features spatially. While conceptually related to our patch-OCR alignment, these approaches operate at pre-training time to improve representations, whereas our approach uses patch similarities at inference time for retrieval filtering. Critically, these models are designed for document understanding tasks (NER, classification) rather than retrieval: they lack late interaction mechanisms and query-conditioned relevance scoring.
OCR-free approaches including Donut (Kim et al., 2022) and Pix2Struct (Lee et al., 2023) perform document understanding directly from pixels. UDoP (Tang et al., 2023) unifies vision, text, and layout modalities through a generative pre-training objective, learning joint representations that could theoretically support retrieval; however, UDoP is optimized for understanding tasks and lacks the late-interaction scoring mechanism essential for efficient large-scale retrieval. LayoutReader (Wang et al., 2023) addresses reading order detection, which affects OCR region quality upstream of our approach. These models excel at understanding but do not address the retrieval problem we target.
@@ 2.3 Region-Level Document Retrieval
RegionRAG. The closest existing work is RegionRAG (Li et al., 2025), which shifts retrieval from document-level to semantic region-level granularity. RegionRAG clusters salient patches via BFS to produce visual region crops, requiring hybrid supervision (bounding box annotations plus pseudo-labels from unlabeled data) and a dual-objective contrastive loss. Our approach differs in three respects: (1) we operate at inference time without additional training, (2) we output text with spatial coordinates rather than image crops, enabling direct use by text-only LLMs, and (3) OCR-derived regions align with semantic document structure (paragraphs, tables, captions) rather than visual patch connectivity. As of December 2025, RegionRAG has no public implementation or model weights available; we therefore compare against the ColPali-family baselines that both approaches share, which represents a limitation of the current comparison landscape rather than a methodological choice.
DocVLM. DocVLM (Shpigel Nacson et al., 2024) integrates OCR into vision-language models by compressing OCR features into learned queries. This represents the opposite direction of our approach. DocVLM adds OCR to enhance VLM understanding, whereas we use late-interaction patch embeddings to filter and score OCR output for retrieval. The positioning of our approach relative to these methods is summarized in Table˜1.
Table 1: Comparison with related approaches. Our approach is unique in achieving region-level retrieval at inference time without additional training, by propagating late-interaction patch similarities to OCR bounding boxes.
@@ 3.1 Problem Formulation
Let $\mathcal{P}=\{p_{1},\ldots,p_{N}\}$ denote the set of $N$ pages in the corpus, where each page $p_{i}$ has an associated set of OCR regions $\mathcal{R}(p_{i})=\{r_{1},\ldots,r_{m}\}$. Each region $r_{j}$ is characterized by its bounding box $B(r_{j})=(x_{1},y_{1},x_{2},y_{2})$ in pixel coordinates and its text content $T(r_{j})$. Our goal is to compute a relevance score $\text{rel}(q,r_{j})$ for each region that captures both semantic relevance and spatial grounding.
@@ 3.2 Coordinate Mapping: Patches to Bounding Boxes
ColPali-family models (including ColQwen3 and ColModernVBERT evaluated in Section˜6) encode each page as a grid of $G\times G$ patch embeddings ($G=32$) over an input image of resolution $I\times I$ ($I=448$). Each patch corresponds to an $s\times s$ pixel region where $s=I/G=14$ pixels. Patches are indexed in raster scan order (left-to-right, top-to-bottom).
@@ 3.3 Relevance Propagation via Patch Similarity
For OCR region $r$ with scaled bounding box $B^{\prime}(r)$, we propagate patch scores via normalized IoU-weighted aggregation:
$\text{rel}(q,r)=\frac{\sum_{j}\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))\cdot\text{score}_{\text{patch}}(j)}{\sum_{j}\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))}$ (6)
where the sums are over all patches $j$ with non-zero intersection. This weights each patch’s contribution by its spatial overlap with the region, while normalizing by total IoU to ensure comparable scores across regions of different sizes.
@@ 3.4 Two-Stage Retrieval Architecture
Stage 1: Candidate Retrieval. We mean-pool the $G^{2}$ patch embeddings to obtain a single $d$-dimensional page-level representation for efficient approximate nearest neighbor search, retrieving top-$K$ candidate pages. This stage uses standard dense retrieval techniques; our contribution focuses on Stage 2.
Stage 2: Region Reranking. For each candidate page, we compute full patch-level similarity and propagate scores to OCR regions as described in Section˜3.3. Regions are ranked by their relevance scores, and top-$k$ regions are returned.
@@ 3.5 Aggregation Strategies
We consider alternative aggregation strategies for propagating patch scores to regions. Let $\text{covered}(r)$ denote the set of patch indices with non-zero intersection with region $r$:
$\text{covered}(r)=\{j:\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))>0\}$ (7)
@@
$\text{rel}_{\text{IoU}}(q,r)=\frac{\sum_{j}\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))\cdot\text{score}_{\text{patch}}(j)}{\sum_{j}\text{IoU}(B^{\prime}(r),\text{patch\_bbox}(j))}$ (10)
This normalizes by total IoU to produce comparable scores across regions of different sizes. Without normalization, larger regions accumulate higher scores simply by covering more patches.
The choice of aggregation strategy affects retrieval quality depending on region size and content density; we use IoU-weighted aggregation as the default based on its principled handling of partial patch overlaps while maintaining size-invariant comparability.
@@ 4 Theoretical Analysis
4.1 Area Efficiency Bounds
The spatial efficiency of our approach is fundamentally bounded by patch resolution. We formalize this tradeoff.
Theorem 1 (Area Efficiency Bound).
For an OCR region with bounding box of width $w$ and height $h$ (in model coordinates), and patch size $s$, the maximum achievable area efficiency (ratio of target region area to total patch area covering it) is:
$\text{efficiency}\leq\frac{w\cdot h}{(w+s)\cdot(h+s)}$ (11)
@@ 4.1 Area Efficiency Bounds
Consider a region with bounding box $B$ of dimensions $w\times h$. The set of patches intersecting $B$ depends on $B$’s alignment with the patch grid. In the worst case, $B$ is positioned such that it intersects partial patches on all four edges. Let the region’s top-left corner fall at position $(x,y)$ within a patch. The region then spans from patch column $\lfloor x/s\rfloor$ to $\lfloor(x+w)/s\rfloor$ and from patch row $\lfloor y/s\rfloor$ to $\lfloor(y+h)/s\rfloor$. The number of intersecting patches is at most $\lceil w/s+1\rceil\cdot\lceil h/s+1\rceil$. The total area covered by these patches is at most $(w+s)\cdot(h+s)$, achieved when the region is maximally misaligned with patch boundaries. The efficiency (ratio of target area to retrieved patch area) is thus bounded by $(w\cdot h)/((w+s)\cdot(h+s))$. ∎
@@
  • A typical paragraph region ($200\times 50$ pixels): efficiency $\leq 73\%$
  • A table cell ($100\times 30$ pixels): efficiency $\leq 60\%$
  • A small label ($50\times 20$ pixels): efficiency $\leq 46\%$
@@ 4.3.1 Context Reduction
4.3.2 Precision and Signal-to-Noise Improvement
Score-based region ranking improves retrieval precision over random selection whenever patch scores correlate positively with relevance, an assumption validated empirically in Section˜6. This precision improvement directly translates to better signal-to-noise ratio in retrieved context: selecting fewer, higher-relevance regions reduces the irrelevant content that downstream LLMs must process. The cost-quality tradeoff in Table˜2 illustrates this directly: our hybrid approach achieves the context efficiency of OCR-based selection while providing the relevance ranking that OCR alone cannot offer.
Table 2: Combined efficiency-quality comparison. The hybrid approach uniquely achieves both low context cost and high precision. Precision values for our approach are empirically measured (IoU@0.5 on BBox-DocVQA). $M\approx 15$ average regions per page.
Method Context Cost Precision (IoU@0.5) Best Use Case